# Shadow-FT: Tuning Instruct via Base

### v2

@@ 5.2 Tuning on Domain Data
Tuning methods are typically employed to adapt LLMs for a specific domain, such as medical. Therefore, we perform tuning experiments on specific domain data, including Medical-o1-reasoning-SFT [41] in the medical domain, Code-Z1 [42] in the code domain, and LIMO [44] & OpenR1-Math  [49] in the math domain. Following LIMO [44], we uniformly down sample the Medical-o1-reasoning-SFT to 1,000, and Code-Z1/OpenR1-Math to 2,000. On these domain tasks, we employ the LoRA with rank 128 and optimize with a learning rate of 2e-4.
Table 2: The detailed mathematical and code performance tuning Qwen-3-8B and Llama-3.1-8B on Code-Z1, LIMO, and OpenR1-Math. Va. denotes the vanilla Instruct baseline, LoRA for conventional LoRA, and Shadow for proposed Shadow-FT (LoRA).
@@
Va. LoRA Shadow LoRA Shadow LoRA Shadow Va. LoRA Shadow LoRA Shadow LoRA Shadow
@@
Figure 4 reports the results of tuning Llama-3.2-1B on Medical-o1-reasoning-SFT. We report the results on MMLU Pro-Medical [47], MedMCQA [46], PubMedQA [45], MMLU-Biology [61], and GPQA-Medical following [41], while normalizing the maximum score to 1 for better visualization. We can find that conventional LoRA would lead to performance degeneration, while Shadow-FT (LoRA) improves the performance, which is consistent with the conclusion on BAAI-2k. Please refer to Appendix B.4 for detailed scores.
Table 2 shows the detailed results of Math-7 and Code-3 tuning Qwen-3-8B and Llama-3.1-8B on Code-Z1, LIMO, and OpenR1-Math. The observations are consistent, i.e., conventional LoRA would lead to degeneration, while the proposed shadow-FT (LoRA) can effectively adapt LLMs on specific domain knowledge. For instance, Shadow-FT (LoRA) achieves a Math-7 score of 77.4 on Qwen-3-8B, which is 6.2 higher than 71.2 of LoRA, and 2.9 higher than the vanilla Instruct model. Moreover, we also find that tuning LLM via Shadow-FT on code data can improve the math capability [42], and vice versa. In particular, when tuned via shadow-FT on Code-z1, the Qwen-3-8B can achieve a score of 36.7 on the tough AIME-24 benchmark, showing superior adaptation and generalization ability.
@@ 5.3 Combined with DPO
Direct Preference Optimization (DPO), which directly optimizes a language model to adhere to human preferences without explicit reward modeling or reinforcement learning, shows promising performance when applying RL to LLMs [25]. Therefore, we try to combine Shadow-FT with DPO, i.e., applying DPO on Base and then grafting the weight to Instruct, termed as Shadow-DPO. Specifically, we achieve Shadow-DPO using LoRA on 1,000 paired samples from the Math-Step [62] dataset and set the rank to 8 and 128. As shown in Table 3, shadow-DPO outperforms DPO under two settings, such as 55.39 vs. 54.62 of vanilla DPO. It shows that the strategy employing the Base as proxy of Instruct also works for DPO. Meanwhile, a larger rank leads to better results for shadow-DPO, which is consistent with results tuning on BAAI-2k shown in Figure 4.
Table 3: Benchmark results when tuning Llama-3.1-8B using DPO and Shadow-DPO. We report Math-7, Code-3, and Reasoning-9 scores with different ranks (8 and 128).
Method Rank Math-7 Code-3 Reasoning-9 Avg.
Vanilla 56.81 50.88 56.61 54.77
DPO 8 57.00 50.26 57.14 54.80
Shadow-DPO 8 57.00 50.61 57.27 54.96
DPO 128 56.24 50.28 57.33 54.62
Shadow-DPO 128 57.48 50.94 57.76 55.39
@@ 5.4 Performance on MLLM
For generality, we further conduct experiments tuning Multimodal Large Language Models (MLLMs). For the dataset, we select 10,000 samples from ChartMoE  [50], which takes a chart and a natural language question as input to predict the answer. For MLLM, we select Gemma-3 [3] 12B/27B and Llama-3.2-Vision [67] 11B/90B. During training, we employ LoRA and set rank as 128. The learning rate is 2e-4. We evaluate the tuned model via lmms-eval framework [68]. As shown in Table 4, both conventional LoRA and Shadow-FT (LoRA) effectively adapt MLLMs on ChartQA [66] tasks. Meanwhile, our proposed Shadow-FT outperforms LoRA especially on larger models, such as 63.80 on Gemma-3-27B compared to 60.28 of vanilla LoRA and 80.6 on Llama-3.2-Vision-90B compared to 79.92.
Table 4: Performance of Gemma-3 and Llama-3.2-Vision on Multi-modal ChartQA tasks.
Model Size Method Aug. Split Human Split Overall
Gemma-3 12B Vanilla 45.20 29.52 37.36
LoRA 61.84 45.12 53.48
Shadow-FT 63.36 46.48 54.92
27B Vanilla 56.56 27.28 41.92
LoRA 70.88 49.68 60.28
Shadow-FT 72.24 55.36 63.80
Llama-3.2-Vision 11B Vanilla 27.04 17.20 22.12
LoRA 93.04 55.84 74.44
Shadow-FT 92.48 55.76 74.12
90B Vanilla 42.56 19.28 30.92
LoRA 93.68 66.16 79.92
Shadow-FT 93.44 67.76 80.60
@@ 5.5 Model Zoo: More LLMs
Model Method Math-7 Code-3 Reasoning-9 Avg.
@@
We further apply Shadow-FT to more LLMs, including Gemma-3 series [3], Yi series [69], and Falcon series [70]. The hyperparameters are the same as tuning Qwen 3 and Llama 3. Table 5 shows the results of Math-7, Code-3, and Reasoning-9. We can find that proposed Shadow-FT consistently outperforms conventional tuning methods. For instance, Shadow-FT gets an average of 52.55 when tuning Gemma-3-4B, which is 1.1 higher than vanilla Instruct model and 7.91 higher than conventional tuning methods. All the tuned models will be made public in the future.
6 Related Work
6.1 Tuning For LLMs
Large language models (LLMs) gain superior ability from pre-training on tremendous data [19], followed by tuning on various downstream tasks [20, 18]. These methods can be categorized into: 1) full-parameters method, which updates all the parameters, and 2) parameter-efficient fine-tuning (PEFT) method, lowering the tuning costs via parameter selection [21] or low-rank branches [14, 22]. More recently, Reinforcement Learning from Human Feedback (RLHF) methods show promising performance in aligning models to human preferences and improving the reasoning ability [25, 2, 23, 24]. These methods focus on improving the training strategy and involve the target model only. In this paper, we propose Shadow-FT to tune Instruct model on Base model. Also, our proposed Shadow-FT can be combined with these baselines to enhance the performance.
6.2 Model Guidance in Tuning
Introducing extra knowledge from other models has been proven as a promising way to enhance tuning performance, such as knowledge distillation [26, 27] and proxy-tuning [8]. Knowledge distillation methods aim to transfer the knowledge from a larger teacher model to a compact student model, via aligning the outputs [27, 28] or employing the teacher’s outputs as training data [29, 30]. Proxy-tuning first tunes a smaller LLM and then applies the logit differences to a larger model [8]. These methods transfer knowledge at the feature level or data level, while our proposed Shadow-FT directly grafts the weight updates. We also notice a very recent concurrent work [32] to transfer the fine-tuning ability. Differently, our proposed Shadow-FT focuses on tuning Instruct via Base model based on the observation that the weights are highly similar. Moreover, we conduct experiments on more LLMs across more benchmarks, and further extend the idea to MLLMs and DPO.
7 Conclusion and Limitations
In Shadow-FT, we first tune the Base model and then graft the weight updates to the Instruct model. However, there are some LLMs for which the paired Base models are not available, such as Qwen-3-32B. For these LLMs, we can not apply Shadow-FT. Therefore, finding a proper ’shadow’ for these models is an interesting topic for future work.
Appendix A Similarity on More LLMs
@@
Appendix B Experimental Details
B.1 Hyper-parameters
@@
B.2 Benchmarks
B.3 Detailed Data of Table 1
@@
B.4 Detailed table for Medical Benchmarks
Table 6 reports the detailed results of tuning Llama-3.2-1B on the Medical-o1-reasoning-SFT dataset.
@@ B.4 Detailed table for Medical Benchmarks
Benchmark Vanilla FT Shadow-FT
GPQA-Medical 23.85 24.10 24.50
MMLU Pro-Medical 25.20 23.95 27.60
MedMCQA 30.15 28.55 32.40
MedQA 25.95 25.60 29.35
PubMedQA 55.85 54.55 60.65
MMLU-Biology 49.68 49.15 51.85
Evaluation Metric Type n-shot CoT

### v3

@@ 5.2 Tuning on Domain Data
Tuning methods are typically employed to adapt LLMs for a specific domain, such as medical. Therefore, we perform tuning experiments on specific domain data, including Medical-o1-reasoning-SFT (Chen et al., 2024) in the medical domain, Code-Z1 (Yu et al., 2025) in the code domain, and LIMO (Ye et al., 2025) & OpenR1-Math (Face, 2025) in the math domain. Following LIMO (Ye et al., 2025), we uniformly down sample the Medical-o1-reasoning-SFT to 1,000, and Code-Z1/OpenR1-Math to 2,000. On these domain tasks, we employ the LoRA with rank 128 and optimize with a learning rate of 2e-4.
[…]
@@
Ins. LoRA Shadow LoRA Shadow LoRA Shadow Ins. LoRA Shadow LoRA Shadow LoRA Shadow
@@
Figure 4 reports the results of tuning Llama-3.2-1B on Medical-o1-reasoning-SFT. We report the results on MMLU Pro-Medical (Wang et al., 2024b), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019), MMLU-Biology (Yue et al., 2024), and GPQA-Medical following (Chen et al., 2024), while normalizing the maximum score to 1 for better visualization. We can find that conventional LoRA would lead to performance degeneration, while Shadow-FT (LoRA) improves the performance, which is consistent with the conclusion on BAAI-2k. Besides, we further report the results directly tuning the Base. Please refer to Appendix D.3 for detailed scores.
Table 2 shows the detailed results of Math-7 and Code-3 tuning Qwen-3-8B and Llama-3.1-8B on Code-Z1, LIMO, and OpenR1-Math. The observations are consistent, i.e., conventional LoRA would lead to degeneration, while the proposed shadow-FT (LoRA) can effectively adapt LLMs on specific domain knowledge. For instance, Shadow-FT (LoRA) achieves a Math-7 score of 77.4 on Qwen-3-8B, which is 6.2 higher than 71.2 of LoRA and 2.9 higher than the vanilla Instruct model. Moreover, we also find that tuning LLM via Shadow-FT on code data can improve the math capability (Yu et al., 2025), and vice versa. In particular, when tuned via shadow-FT on Code-z1, the Qwen-3-8B can achieve a score of 36.7 on the tough AIME-24 benchmark, showing superior adaptation and generalization ability.
5.3 Mechanistic Analysis of Optimization Dynamics
Figure 5: Optimization dynamics on loss and gradient when tuning Qwen3-8B via Instruct (i.e., FT) and Base (i.e., Shadow-FT).
To provide insight into why Shadow-FT outperforms vanilla FT, we further analyze the optimization dynamics of both methods from a loss and gradient perspective. We denote the loss and gradient for the INSTRUCT model (tuned with vanilla FT) as Loss(I) and Grad(I), and for the BASE model (tuned with Shadow-FT) as Loss(S) and Grad(S).
[…]
Overall, the figure reveals a fundamental contrast in optimization. Vanilla FT on the Instruct model is hindered by a large initial gradient that rapidly diminishes, while the Base model can avoid this and shows a stable trajectory. However, Base model is a good learner but a poor backbone due to the lack of post-training. The conclusion is consistent with Pass@k results detailed in Appendix D.4. Therefore, we propose Shadow-FT to learn on Base and execute on Instruct.
@@ 5.4 Combined with DPO
Direct Preference Optimization (DPO), which directly optimizes a language model to adhere to human preferences without explicit reward modeling or reinforcement learning, shows promising performance when applying RL to LLMs (Rafailov et al., 2023). Therefore, we try to combine Shadow-FT with DPO, i.e., applying DPO on Base and then grafting the weight to Instruct, termed as Shadow-DPO. Specifically, we achieve Shadow-DPO using LoRA on 1,000 paired samples from the Math-Step (Lai et al., 2024) dataset and set the rank to 8 and 128. As shown in Table 5.5, shadow-DPO outperforms DPO under two settings, such as 55.39 vs. 54.62 of vanilla DPO. It shows that the strategy employing the Base as proxy of Instruct also works for DPO. Meanwhile, a larger rank leads to better results for shadow-DPO, which is consistent with results tuning on BAAI-2k shown in Figure 4.
@@ 5.5 Performance on MLLM
For generality, we further conduct experiments tuning Multimodal Large Language Models (MLLMs). For the dataset, we select 10,000 samples from ChartMoE  (Xu et al., 2024), which takes a chart and a natural language question as input to predict the answer. For MLLM, we select Gemma-3 (Team et al., 2025a) 12B/27B and Llama-3.2-Vision (Grattafiori et al., 2024) 11B/90B. During training, we employ LoRA and set the rank to 128. The learning rate is 2e-4. We evaluate the tuned model via lmms-eval framework (Zhang et al., 2024). As shown in Table 5.5, both conventional LoRA and Shadow-FT (LoRA) effectively adapt MLLMs on ChartQA (Masry et al., 2022) tasks. Meanwhile, our proposed Shadow-FT outperforms LoRA, especially on larger models, such as 63.80 on Gemma-3-27B compared to 60.28 of vanilla LoRA and 80.6 on Llama-3.2-Vision-90B compared to 79.92.
Table 3: Performance of tuning Llama-3.1-8B using DPO and Shadow-DPO with ranks be 8 and 128, respectively.
Method Rank 19-Avg.
Vanilla 54.77
DPO 8 54.80
Shadow-DPO 8 54.96
DPO 128 54.62
Shadow-DPO 128 55.39
Table 4: Performance of Gemma-3 and Llama-3.2-Vision on the multi-modal ChartQA task. We set the rank of LoRA to 128.
Model Size Vanilla LoRA Shadow-FT
Gemma-3 12B 37.36 53.48 54.92
27B 41.92 60.28 63.80
Llama-3.2 11B 22.12 74.44 74.12
-Vision 90B 30.92 79.92 80.60
5.6 Weight Delta Scaling
In Shadow-FT, we directly graft the learned weights from BASE to Instruct. We further explore the scaling of learned weights. Please refer to Appendix D.6 for more details. In summary, our proposed Shadow-FT outperforms vanilla Instruct with different scaling factors, showing strong robustness. Moreover, a factor slightly larger than 1 would yield better results, while we leave it for future work to explore the best factor.
@@ 5.7 Model Zoo: More LLMs
We further apply Shadow-FT to more LLMs, including Gemma-3 series (Team et al., 2025a), Yi series (Young et al., 2024), and Falcon series (Almazrouei et al., 2023). Please refer to Appendix B for more details. We can find that proposed Shadow-FT consistently outperforms conventional tuning methods. All the tuned models will be made public in the future.
6 Related Work
6.1 Tuning For LLMs
Large language models (LLMs) gain superior ability from pre-training on tremendous data (Gururangan et al., 2020), followed by tuning on various downstream tasks (Ouyang et al., 2022; Muennighoff et al., 2025). These methods can be categorized into: 1) full-parameters method, which updates all the parameters, and 2) parameter-efficient fine-tuning (PEFT) method, lowering the tuning costs via parameter selection (Zaken et al., 2021) or low-rank branches (Hu et al., 2022; Wu et al., 2024b). More recently, Reinforcement Learning from Human Feedback (RLHF) methods show promising performance in aligning models to human preferences and improving the reasoning ability Rafailov et al. (2023); Bai et al. (2023); Guo et al. (2025); Team et al. (2025b). These methods focus on improving the training strategy and involve the target model only. In this paper, we propose Shadow-FT to tune Instruct model on Base model. Also, our proposed Shadow-FT can be combined with these baselines to enhance the performance.
6.2 Model Guidance in Tuning
Introducing extra knowledge from other models has been proven as a promising way to enhance tuning performance, such as knowledge distillation (Hinton et al., 2015; Wu et al., 2024a) and proxy-tuning (Liu et al., 2024). Knowledge distillation methods aim to transfer the knowledge from a larger teacher model to a compact student model, via aligning the outputs (Wu et al., 2024a; Yang et al., 2024b) or employing the teacher’s outputs as training data (Qin et al., 2024; Min et al., 2024). Proxy-tuning first tunes a smaller LLM and then applies the logit differences to a larger model (Liu et al., 2024). These methods transfer knowledge at the feature level or data level, while our proposed Shadow-FT directly grafts the weight updates. RE-Adapt (Fleshman & Durme, 2024) also utilizes the Base/Instruct model pair for adaptation. However, RE-Adapt models the static weight difference with a low-rank approximation, whereas Shadow-FT is a model-free approach that directly transfers the full dynamic updates without any assumption. Additionally, we notice a very recent concurrent work (Lin et al., 2025) to transfer the fine-tuning ability. Differently, our proposed Shadow-FT focuses on tuning Instruct via Base model based on the observation that the weights are highly similar. Moreover, we conduct experiments on more LLMs across more benchmarks, and further extend the idea to MLLMs and DPO.
7 Conclusion
Reproducibility Statement
We are committed to the reproducibility of our work. The MATH-7 results, averaged across three trials, are shown in Table 1 and Table 8. The full source code required to reproduce our experiments is included in the supplementary material. Corresponding hyperparameters and detailed configuration files for all experiments are documented in Section 4.1. All experiments were conducted on publicly available benchmarks, and the details are provided in Appendix A.
Ethics statement
We have adhered to the ICLR Code of Ethics in this research. Our work is based entirely on publicly available models and benchmarks, involves no human subjects, and we commit to releasing our code to ensure reproducibility.
Appendix A Benchmarks Details
Evaluation Metric Type n-shot CoT
Math-7
AIME24 pass@1 sampling 0-shot
GSM8K(0-shot) Accuracy sampling 0-shot
GSM8K(8-shot) Accuracy sampling 8-shot
MATH Accuracy sampling 0-shot
MATH-500 Accuracy sampling 0-shot
Minerva Math Accuracy sampling 4-shot
SVAMP Accuracy sampling 0-shot
Code-3
HumanEval pass@1 sampling 0-shot
HumanEval+ pass@1 sampling 0-shot
LiveCodeBench average
- generation pass@1 sampling 0-shot
- test pass@1 sampling 0-shot
- prediction pass@1 sampling 0-shot
Knowledge-9
ARC-Challenge Accuracy sampling 0-shot
BBH(0-shot) Accuracy sampling 0-shot
BBH(3-shot) Accuracy sampling 3-shot
Drop Accuracy sampling 0-shot
GPQA Diamond Accuracy sampling 0-shot
MMLU Accuracy sampling 0-shot
MMLU Pro Accuracy sampling 0-shot
Winogrande Accuracy sampling 0-shot
TheoremQA Accuracy sampling 0-shot
Appendix B Model Zoo: More LLMs
@@
Model Method Math-7 Code-3 Knowledge-9 Avg.
@@
We further apply Shadow-FT to more LLMs, including Gemma-3 series (Team et al., 2025a), Yi series (Young et al., 2024), and Falcon series (Almazrouei et al., 2023). The hyperparameters are the same as tuning Qwen 3 and Llama 3. Table 6 shows the results of Math-7, Code-3, and Knowledge-9. We can find that proposed Shadow-FT consistently outperforms conventional tuning methods. For instance, Shadow-FT gets an average of 52.55 when tuning Gemma-3-4B, which is 1.1 higher than the vanilla Instruct model and 7.91 higher than conventional tuning methods.
Appendix C Similarity on More LLMs
@@
Appendix D Experimental Details
D.1 Hyper-parameters
@@
D.2 Detailed Data of Table 1
@@
D.3 Detailed table for Medical Benchmarks
Table 7 reports the detailed results of tuning Llama-3.2-1B on the Medical-o1-reasoning-SFT dataset. Shadow-FTR donates the method integrating fine-tuned weights from INSTRUCT to BASE.
@@ D.3 Detailed table for Medical Benchmarks
Benchmark Tune on Instruct Tune on Base
Instruct FT Shadow-FT Base Base-FT Shadow-FTR
GPQA-Medical 23.85 24.10 24.50 25.25 25.00 24.75
MMLU Pro-Medical 25.20 23.95 27.60 13.10 12.60 12.30
MedMCQA 30.15 28.55 32.40 30.20 30.60 29.50
MedQA 25.95 25.60 29.35 29.80 30.35 29.45
PubMedQA 55.85 54.55 60.65 49.35 51.90 50.20
Avg. 32.20 31.35 34.90 29.54 30.09 29.24
D.4 Performance on Pass@k
To evaluate the exploration capability, we use the popular Pass@k, which is defined as the fraction of problems for which at least one correct response is produced in $k$ independent trials. However, directly computing Pass@k using only $k$ rollouts for each problem often suffers from high variance. Therefore, we adapt the unbiased estimator (Chen et al., 2021b). Specifically, we roll out for $n$ times ($n\geq k$), and calculate Pass@k as follows:
$\text{Pass}@k:=\mathbb{E}_{x\sim\mathcal{D}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$], (5)
where $x$ is the input prompt from dataset $D$, and $c$ is the count of correct solutions.
Figure 7: Pass@k performance on Qwen3-8B thinking and non-thinking modes.
We employ the Qwen3-8B model, which supports seamless switching between Thinking mode and Non-thinking mode. We can easily switch between two modes using one control hyperparameter. We set other hyperparameters following the official report of Qwen3 (Team, 2025): do_sample=True, temperature=0.6, top_k=20, top_p=0.95, max_new_tokens=38912 for bettr alignment.
Non-thinking mode.
Under non-thinking decoding, absolute Pass@k values are small for all methods, yet Shadow-FT exhibits a clearer upward trend with larger $k$, progressively surpassing the Instruct baseline. In contrast, vanilla FT yields weak performance and rarely produces correct solutions. Importantly, all methods used the same number of training examples and the same training cutoff length; the only difference is the initialization (Base vs. Instruct). This comparison suggests that Base is a better learner for supervised adaptation—its newly acquired knowledge is less prone to collapse.
Thinking mode.
Under thinking mode, the effect is more pronounced: Shadow-FT and Instruct follow similarly steep, rapidly saturating curves that nearly reach the model’s capacity limit, whereas vanilla FT maintains a large (>>30%) gap across $k$. This pattern implies that vanilla FT hurts the thinking ability of Instruct and makes it less receptive to new knowledge compared with Base-initialized training.
Across both modes, our proposed Shadow-FT avoids collapse and retains—often enhances the upper-bound competence of the underlying Instruct model. This property is valuable for subsequent RL or other generalization-critical settings (Zhu et al., 2025). We attribute the robustness to the favorable inductive characteristics of Base-initialized learning, whereas vanilla FT on an Instruct model struggles to achieve the same balance of stability and adaptability.
D.5 Ranks in LoRA on Llama-3.1-8B
Figure 8: Performance tuning Llama-3.1-8B with different ranks.
D.6 Weight Delta Scaling
To further explore the effect of scaling the transferred deltas in the Shadow-FT strategy, we introduce an interpolation design controlled by a scaling factor $\alpha$. Specifically, let $\Delta W_{B}=W_{B}^{+}-W_{B}$ denote the weight updates learned on the Base model. Instead of directly applying the full deltas to the Instruct model, we interpolate as follows:
$W_{I}^{+}=W_{I}+\alpha\cdot\Delta W_{B}=W_{I}+\alpha\cdot(W_{B}^{+}-W_{B}).$ (6)
Here, $\alpha=1.0$ recovers the standard Shadow-FT formulation, while $\alpha=0.0$ reduces to the original Instruct model without transfer. Intermediate values of $\alpha$ provide a smooth interpolation between the two, allowing us to examine how the magnitude of transferred deltas affects downstream performance. The summarized results of Llama-3.1-8B-Instruct on the MATH-7 benchmark are presented in Table 8, showing that $\alpha=1.0$ yields the excellent overall trade-off, while other ratios offer insights into the sensitivity of tasks to partial transfer. Adaptive scaling strategies (e.g., layer-wise or task-specific factors) are left for future work.
Table 8: Detailed results on the math benchmarks (averaged over three repeated runs for each ratio). The ratio of 0.0 denotes vanilla Instruct, while 1.0 for the proposed Shadow-FT.
Ratio Math-7 AIME24
GSM8K
(8-shot)
GSM8K
(0-shot)
MATH
MATH
500
Minerva
Math
SVAMP
0.0 56.2 1.1 84.8 85.0 48.5 50.7 40.5 82.7
0.1 56.5 1.1 85.2 84.3 49.8 51.1 41.3 83.0
0.2 57.1 2.2 85.0 85.0 50.4 50.5 42.3 84.3
0.3 57.9 2.2 85.0 84.4 51.3 53.5 43.8 85.0
0.4 58.8 6.7 85.6 83.6 52.9 54.2 44.2 84.7
0.5 59.0 6.7 86.0 83.7 52.8 53.7 45.6 84.7
0.6 59.1 4.4 86.1 83.5 52.8 54.7 47.2 84.7
0.7 59.6 5.6 85.1 84.0 53.6 55.8 47.6 85.3
0.8 59.4 5.6 84.9 84.5 53.1 53.6 48.8 85.3
0.9 59.8 3.3 85.0 84.7 53.0 56.0 49.4 87.0
1.0 59.5 3.3 84.7 85.1 53.0 54.9 49.4 85.8
1.2 59.1 0.0 84.8 86.0 52.6 56.6 49.1 84.3
1.5 59.9 3.3 84.6 84.9 52.4 55.9 50.5 87.3
2.0 57.9 0.0 83.2 84.2 50.2 52.0 48.1 87.7
Table 9: Detailed results on the math benchmarks for Table 1. Three times independent training and three times evaluation averages are reported.
Model Method AIME24
GSM8K
(8-shot)
GSM8K
(0-shot)
MATH
MATH
500
Minerva
Math
SVAMP